# CTNeRF: Cross-Time Transformer for Dynamic Neural Radiance Field from Monocular Video

### v1

@@ III-B1 Cross-Time Aggregation
III-B2 Ray Aggregation
=Ray $\mathcal{F}=\text{Ray\ Transformer}(\hat{F}_{t\pm i})$ (15)
Specifically, we define:
$\mathcal{F}=Softmax(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}})\cdot V$ (17)
III-C Frequency Domain Aggregation Module
We introduce a novel spatio-temporal feature learning module termed the Global Spatio-Temporal Filter (GSTF), inspired by recent advancements in frequency-domain-based methodologies aimed at enhancing the rendering quality of new views [69]. The primary objective of GSTF is to elevate the representation of feature vectors by capturing both spatial and temporal relationships through specialized frequency filters. In our approach, GSTF is meticulously crafted to discern and learn distinct frequency filters at each spatial location. This enables the modeling of temporal variations within feature vectors across different spatial positions. The core mechanism involves the transformation of both temporal and spatial features at each location into frequency feature spectra. This transformation is achieved through a two-dimensional Fast Fourier Transform (FFT) [70]. The frequency filter, learned through GSTF, acts as a modulator on this transformed spectrum. Subsequently, we revert this modulated spectrum back to the time domain using an inverse FFT. This comprehensive process allows GSTF to effectively encode the intricate interplay between time and space in the feature vectors, contributing to the improvement of new view rendering. To gain a better understanding of our GSTF design, let’s first review the convolution theorem in the field of digital signal processing [71]. Given a sequence of feature signals with $T$ points ($f[t],0\leq t\leq T-1$), we can calculate its discrete spectrum $S[k]$ using Discrete Fourier Transform (DFT) via:
$S[k]=\sum_{t=0}^{T-1}f[t]e^{-j(2\pi/T)kt},0\leq k\leq T-1$ (18)
$f[t]=\frac{1}{T}\sum_{k=0}^{T-1}S[k]e^{j(2\pi/T)kt},0\leq t\leq T-1$ (19)
Our GSTF can be easily used in modern deep-learning frameworks such as PyTorch [72]. Through ablation experiments, we demonstrate that our global filtering mechanism is an effective spatial information mixing method. The GSTF module improves the texture quality of new views and alleviates blurring and artifacts often observed in new view synthesis.
Figure 6: Network architecture of our static representation.
Figure 7: Network architecture of our dynamic representation.
III-D Regularization
$\mathcal{L}=\mathcal{L}_{data}+\mathcal{L}_{small}+\mathcal{L}_{pho}$ (20)
IV Experiment
IV-A Implementation details
IV-B Architecture
Our approach utilizes two distinct types of models: static and dynamic. The architecture of the static model is depicted in Figure 6, while the architecture of the dynamic model is illustrated in Figure 7.
IV-C Dataset
Nvidia Dynamic Scene Dataset comprises nine video sequences captured using a static camera rig of 12 cameras. All cameras capture images simultaneously at 12 different time steps $\{t_{0},t_{1},\dots,t_{11}\}$, and we obtain a twelve-frame monocular video $\{I_{0},I_{1},\dots,I_{11}\}$ by sampling the image taken by the $i-th$ camera at time $t_{i}$. It is worth mentioning that we use a different camera for each video frame to simulate camera motion. The video frames consist of a background that remains stationary throughout the video and a dynamic object that changes over time. We adopt COLMAP [74] similar to NeRF [11], to estimate the camera poses, near and far boundaries of the scene, and assume that all cameras share the same intrinsic parameter. We exclude the DynamicFace sequence from our evaluation since COLMAP fails to estimate camera poses for this sequence. Lastly, we resize all video sequences to a resolution of 480 $\times$ 270.
The DAVIS Dataset [62] consists of fifty sequences featuring dynamic moving objects, like animals and cars. However, due to limitations in camera movement, COLMAP could only estimate camera poses for six out of the fifty sequences, all of which include ground truth object masks.
Finally, for the iPhone Dataset [60], we conducted a quantitative evaluation using seven dynamic scenes, each accompanied by ground truth images of novel views.
IV-D Comparison with State-of-the-Art Methods
IV-D1 Qualitative Results
IV-D2 Quantitative Results
Table I presents the quantitative results obtained from the Nvidia Dynamic Scene Dataset [58]. We adopted the evaluation methodology from DynamicNeRF [16] to synthesize views using the first camera while varying the time on the Nvidia Dynamic Scene Dataset. To evaluate the rendering quality of each method, we employed two widely recognized error metrics: peak signal-to-noise ratio (PSNR) and perceptual similarity (LPIPS) as defined by [65]. Additionally, due to minor differences observed in our ablation study’s results, we incorporated the structural similarity index (SSIM) for a more thorough assessment.
Our method demonstrates significant advancements, outperforming existing state-of-the-art techniques in five of the seven tested scenarios. This improvement is particularly evident in the average PSNR increase of 1dB and a notable 20% reduction in LPIPS error, underscoring a substantial enhancement in perceptual quality compared to real images. Moreover, we extended our evaluation following DyCheck’s methodology [60] for the iPhone dataset, detailed in Table Table II, where we report masked PSNR and SSIM scores. Given that a significant number of scenes within this dataset are long sequences, and considering our method’s limitations in effectively modeling such sequences, notable improvements are limited. Nevertheless, our method demonstrates comparable performance to established methods and even exhibits slight enhancements in select scenarios. These outcomes serve as a compelling demonstration of the superior effectiveness of our framework in restoring intricate scene content.
@@ IV-D2 Quantitative Results
TABLE III: Evaluate the whole module on the Nvidia Dynamic Scene Dataset [58] (Balloon 2 scene).
@@
IV-E Ablation Study
IV-E1 Evaluate the whole module
In Table III, we present a detailed comparison between our complete system and its variants, each lacking a specific module: A) multi-view aggregation, B) an additional four-layer MLP, C) Global Spatio-Temporal Filter module, D) Ray-Based Cross-Time Aggregation Module, and E) regularization scene flow loss. As indicated in the first two rows of Table III, the absence of multi-view aggregation and the additional four-layer MLP markedly diminish the quality of view synthesis, with a decrease in PSNR by 0.75% and 2.61% respectively, and SSIM by 1.29% and 2.81% respectively. Additionally, LPIPS scores increased by 25.00% and 27.08%, signaling a noticeable degradation in image quality and accuracy. These two components, therefore, are critical in enhancing the fidelity and precision of the synthesized views. We observe that removing the Global Spatio-Temporal Filter module and the Ray-Based Cross-Time Aggregation Module also impacts the system’s performance, though to a slightly lesser extent compared to the first two components, with PSNR decreasing by 2.08% and 0.63%, and SSIM by 2.23% and 0.44% respectively. Additionally, the removal of the regularization scene flow loss demonstrates a relatively minor impact on the quality of view synthesis, with a less pronounced decrease in performance metrics compared to the other modules. This suggests that while this component aids in fine-tuning the system, its absence does not drastically compromise the overall effectiveness.
TABLE IV: Evaluate the RBCT module on the Nvidia Dynamic Scene Dataset [58] (Balloon 2 scene).
@@ IV-E1 Evaluate the whole module
TABLE V: Evaluate the temporal module on the Nvidia Dynamic Scene Dataset [58] (Balloon 2 scene).
@@
IV-E2 Evaluate the RBCT module
@@
Furthermore, we investigate the impact of the internal structure of the RBCT module on the model-view synthesis performance, as summarized in Table IV. Specifically, we examine the effects of the following variations: A) without using Cross-time Transformer, B) without using Ray Transformer, C) without using global ray sampling point coordinate embedding, and D) using Ray Transformer first, followed by Cross-time Transformer. Excluding the Cross-time Transformer led to a moderate decline in synthesis quality, as indicated by a 0.74% decrease in PSNR and a 1.29% drop in SSIM, coupled with a notable 25% increase in LPIPS. The omission of the Ray Transformer had a more pronounced impact on performance, with a 2.61% reduction in PSNR, a 2.82% decrease in SSIM, and a 27.08% rise in LPIPS. This highlights the Ray Transformer’s critical role in maintaining high-quality synthesis. Furthermore, removing the global ray sampling point coordinate embedding also negatively affected the results, leading to a 2.08% reduction in PSNR, a 2.23% decrease in SSIM, and a 16.67% increase in LPIPS. Sequentially applying the Ray Transformer and the Cross-time Transformer slightly improved some metrics compared to using the full module configuration, with a 0.11% increase in PSNR and a 0.53% rise in SSIM, while maintaining stable LPIPS. Therefore, our experiments demonstrated that the absence of key components, especially the Ray Transformer and the global ray sampling point coordinate embedding, significantly compromises view synthesis quality. In Figure 8, the absence of global ray sampling point coordinate embedding resulted in stripes in the dynamic synthesis region. Although the other two experiments have a minimal impact on the model, the differences are discernible.
IV-E3 Evaluate the temporal module
To demonstrate the effectiveness of our proposed frequency-domain timing module, as summarized in Table V. Specifically, A) without using Cross-time Transformer, B) Global Spatio-Temporal Filter module. The model without CTT shows a 1.11% improvement in PSNR, a 0.41% improvement in SSIM, and a 33.33% reduction in LPIPS compared to the full model. Conversely, without GSTF, although the PSNR improves by 0.1%, SSIM increases by 2.34%, and LPIPS decreases by 16.67%. In the Full model, we observe more substantial improvements. Compared to A), the Full model shows a 1.2% increase in PSNR, a 1.11% improvement in SSIM, and a 20.00% reduction in LPIPS. In comparison to B), the Full model exhibits a 0.1% increase in PSNR, a 0.23% decrease in SSIM, and a 4.00% reduction in LPIPS. This indicates that simultaneous utilization of CTT and GSTF significantly enhances the quality of the novel view, with a more pronounced improvement in perceptual quality. From the perspective of the error metric, there may not be a significant disparity between the two approaches, but utilizing them simultaneously can be complementary and enhance the quality of the novel view.
TABLE VI: Evaluation of GSTF module for Dynamic region on the Nvidia Dynamic Scene Dataset [58] (Balloon 2 scene).
@@ IV-E3 Evaluate the temporal module
IV-E4 Evaluate the GSTF module
In Figure 10, we present a visual comparison between the baseline features extracted from Resnet34 and the features refined by our GSTF module. Our GSTF is designed with the specific goal of capturing detailed contours and high-frequency texture information, ensuring the preservation of sharp textures in the reconstructed view. A quantitative evaluation in Table VI further underscores the impact of the GSTF module, particularly in the context of dynamic scenes. The results reveal a substantial 2.1% increase in PSNR, a noteworthy 4.68% improvement in SSIM, and a significant 25.38% reduction in LPIPS when utilizing the GSTF module (w/ GSTF) compared to the configuration without GSTF (w/o GSTF). It’s worth highlighting that in Table III, the GSTF module makes a marginal contribution to the overall improvement. This is primarily because our GSTF is applied to dynamic areas, which represent just one-fifth of the total in the Balloon 2 evaluation scene.
IV-E5 Ablation study qualitative results
@@ IV-F Limitations
V Conclusion
In this work, we present a novel dynamic neural rendering field framework for dynamic monocular videos, which allows for the high-quality rendering of novel views. To achieve this, we have extended the recent idea of multi-view aggregation to time-varying NeRF, which enables the modeling of complex motion. Specifically, we introduce the RBCT and the GSTF modules to model motion from the time and frequency domains, respectively. Our experimental results demonstrate that these proposed modules significantly enhance the performance of time-varying NeRF with multi-view aggregation when rendering novel views.
While our work represents a promising exploration of time-varying NeRF for multi-view aggregation, there are still some limitations. Notably, our current method may not perform well when rendering novel views for long-sequence videos. One potential solution to improve performance is to increase the length of the aggregation view, but this approach would require significant computing resources.

### v2

@@ 4.4.2 Quantitative Results
Table 4: Evaluation of the whole module, RBCT module, and temporal module on the Nvidia Dynamic Scene Dataset [44] (Balloon 2 scene).
@@
(a) Whole Module
@@
(b) RBCT Module
@@
(c) Temporal Module
Our method demonstrates significant advancements, outperforming existing state-of-the-art techniques in five of the seven tested scenarios. This improvement is particularly evident in the average PSNR increase of 1dB and a notable 20% reduction in LPIPS error, underscoring a substantial enhancement in perceptual quality compared to real images. Moreover, we extended our evaluation following DyCheck’s methodology [31] for the iPhone dataset, detailed in Table Table 3, where we report masked PSNR and SSIM scores. Given that a significant number of scenes within this dataset are long sequences, and considering our method’s limitations in effectively modeling such sequences, notable improvements are limited. Nevertheless, our method demonstrates comparable performance to established methods and even exhibits slight enhancements in select scenarios. These outcomes serve as a compelling demonstration of the superior effectiveness of our framework in restoring intricate scene content.
4.5 Ablation Study
4.5.1 Evaluate the whole module
In Table 4, we present a detailed comparison between our complete system and its variants, each lacking a specific module: A) multi-view aggregation, B) an additional four-layer MLP, C) Global Spatio-Temporal Filter module, D) Ray-Based Cross-Time Aggregation Module, and E) regularization scene flow loss. As indicated in the first two rows of Table 4, the absence of multi-view aggregation and the additional four-layer MLP markedly diminish the quality of view synthesis, with a decrease in PSNR by 0.75% and 2.61% respectively, and SSIM by 1.29% and 2.81% respectively. Additionally, LPIPS scores increased by 25.00% and 27.08%, signaling a noticeable degradation in image quality and accuracy. These two components, therefore, are critical in enhancing the fidelity and precision of the synthesized views. We observe that removing the Global Spatio-Temporal Filter module and the Ray-Based Cross-Time Aggregation Module also impacts the system’s performance, though to a slightly lesser extent compared to the first two components, with PSNR decreasing by 2.08% and 0.63%, and SSIM by 2.23% and 0.44% respectively. Additionally, the removal of the regularization scene flow loss demonstrates a relatively minor impact on the quality of view synthesis, with a less pronounced decrease in performance metrics compared to the other modules. This suggests that while this component aids in fine-tuning the system, its absence does not drastically compromise the overall effectiveness.
4.5.2 Evaluate the RBCT module
@@
Furthermore, we investigate the impact of the internal structure of the RBCT module on the model-view synthesis performance, as summarized in Table 4. Specifically, we examine the effects of the following variations: A) without using Cross-time Transformer, B) without using Ray Transformer, C) without using global ray sampling point coordinate embedding, and D) using Ray Transformer first, followed by Cross-time Transformer. Excluding the Cross-time Transformer led to a moderate decline in synthesis quality, as indicated by a 0.74% decrease in PSNR and a 1.29% drop in SSIM, coupled with a notable 25% increase in LPIPS. The omission of the Ray Transformer had a more pronounced impact on performance, with a 2.61% reduction in PSNR, a 2.82% decrease in SSIM, and a 27.08% rise in LPIPS. This highlights the Ray Transformer’s critical role in maintaining high-quality synthesis. Furthermore, removing the global ray sampling point coordinate embedding also negatively affected the results, leading to a 2.08% reduction in PSNR, a 2.23% decrease in SSIM, and a 16.67% increase in LPIPS. Sequentially applying the Ray Transformer and the Cross-time Transformer slightly improved some metrics compared to using the full module configuration, with a 0.11% increase in PSNR and a 0.53% rise in SSIM, while maintaining stable LPIPS. Therefore, our experiments demonstrated that the absence of key components, especially the Ray Transformer and the global ray sampling point coordinate embedding, significantly compromises view synthesis quality. In Figure 7, the absence of global ray sampling point coordinate embedding resulted in stripes in the dynamic synthesis region. Although the other two experiments have a minimal impact on the model, the differences are discernible.
4.5.3 Evaluate the temporal module
@@
Table 5: Evaluation of GSTF module for Dynamic region on the Nvidia Dynamic Scene Dataset [44] (Balloon 2 scene).
To demonstrate the effectiveness of our proposed frequency-domain timing module, as summarized in Table 4. Specifically, A) without using Cross-time Transformer, B) Global Spatio-Temporal Filter module. The model without CTT shows a 1.11% improvement in PSNR, a 0.41% improvement in SSIM, and a 33.33% reduction in LPIPS compared to the full model. Conversely, without GSTF, although the PSNR improves by 0.1%, SSIM increases by 2.34%, and LPIPS decreases by 16.67%. In the Full model, we observe more substantial improvements. Compared to A), the Full model shows a 1.2% increase in PSNR, a 1.11% improvement in SSIM, and a 20.00% reduction in LPIPS. In comparison to B), the Full model exhibits a 0.1% increase in PSNR, a 0.23% decrease in SSIM, and a 4.00% reduction in LPIPS. This indicates that simultaneous utilization of CTT and GSTF significantly enhances the quality of the novel view, with a more pronounced improvement in perceptual quality. From the perspective of the error metric, there may not be a significant disparity between the two approaches, but utilizing them simultaneously can be complementary and enhance quality of novel view.
4.5.4 Evaluate the GSTF module
In Figure 9, we present a visual comparison between baseline features extracted from Resnet34 and the features refined by our GSTF module. Our GSTF is designed with the specific goal of capturing detailed contours and high-frequency texture information, ensuring the preservation of sharp textures in the reconstructed view. A quantitative evaluation in Table 5 further underscores the impact of the GSTF module, particularly in the context of dynamic scenes. The results reveal a substantial 2.1% increase in PSNR, a noteworthy 4.68% improvement in SSIM, and a significant 25.38% reduction in LPIPS when utilizing the GSTF module (w/ GSTF) compared to the configuration without GSTF (w/o GSTF). In Table 4, the GSTF module makes a marginal contribution to the overall improvement. This is primarily because our GSTF is applied to dynamic areas, which represent just one-fifth of the total in the Balloon 2 evaluation scene.
4.5.5 Ablation study qualitative results
@@ 4.6 Limitations
5 Conclusion
In this work, we aim to introduce a novel dynamic neural render field framework for dynamic monocular videos, which enables high-quality rendering of novel views. To achieve this goal, we extend recent ideas in multi-view aggregation to time-varying NeRF, enabling the modeling of complex motion. Specifically, we introduce RBCT and GSTF modules to model motion from the time domain and frequency domain, respectively. Our experimental results show that these proposed modules significantly improve the performance of time-varying NeRF with multi-view aggregation when rendering new views. While our work represents a promising exploration of time-varying NeRF for multi-view aggregation, there are still some limitations. It is worth noting that our current method may not perform well when rendering novel views of long sequences of videos. One potential solution to improve performance is to increase the length of the aggregate view, but this approach requires significant computing resources. Fortunately, recent developments such as TensoRF and 3D Gaussian splatting offer potential solutions to these challenges.